OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 2x^2 - 2x = 1

OpenStudy (anonymous):

Quadratic formula is b^2 - 4ac where a,b,c are coefficients of the quadratic equation a.x^2 + b.x + c = 0 Therefore, in this case, a is 2 and b is -2 and c is -1. Hence, The solution is (-2)^2 - 4(2)(-1) = 12 Hope this helps.

OpenStudy (anonymous):

2x^2-2x-1=0 \[x=(-b \pm \sqrt{b^2-4ac})/2a\] now a=2,b=-2,c=-1 \[x=(2\pm \sqrt{4+8})/4=2\pm \sqrt{12}/4=2\pm2\sqrt{3}/4=1\pm \sqrt{3}/2\]

OpenStudy (anonymous):

\[x=(1\pm \sqrt{3})/2\]